A rectangular garden is surrounded by a walk of uniform width. The area of the garden is 180 square yards. If the dimentions of the garden plus the walk are 16 yards by 25 yards, find the width of the walk.
Answer by mananth(16946)

A rectangular garden is surrounded by a walk of uniform width. The area of the garden is 180 square yards. If the dimentions of the garden plus the walk are 16 yards by 25 yards, find the width of the walk.
...
let the width of the sidewalk be x
dimensions of garden will be
25-2x , 16-2x
Area = L*W
..
(25-2x)(16-2x)=180
400-50x-32x+4x^2=180
4x^2-82x+400-180=0
4x^2-82x+220=0
/2
2x^2-41x+110=0
discriminant = 801
x1= (41+sqrt(801))/4
x1=17.32 yards ( not possible )
x2=(41-sqrt(801))/4
x2= 3.17 yards
...
m.ananth@hotmail.ca